1. Assume that you are given values in eax, ebx, ecx. Write an assembly code that does the following: eax = (ecx + edx ) - (eax + ebx)

2. Write a piece of code that copies the number inside al to ch. Example: Assume that Initially eax = 0x15DBCB19. At the end of your code ecx = 0x00001900. Your code must be as efficient as possible.

3. You are given eax = 0x5. Write one line of code in assembly to Zero eax. Don't use Sub instruction and mov instruction. You must use hex numbers only as immediate numbers.

4. Implement the following arithmetic expression in assembly language: EBX = var1 - var2 - 9 + var3 // size of variables, 4 bytes each

5. Write a piece of code in assembly that implements the following C++ statement num1 = num2 = 4; // x and y are of size word (2 bytes)

6. Write a piece of code that implements the following C++ statements d = 'w'; // a ,d,b are of size byte. a = d + 1; b = a + 2;

Part two:(8 points) Open hello.asm. Before the exit line of code, add the line of code: mov ebx, -20 add ebx, 1 Save and build the project. Set a breakpoint after the add line that you just added. Run to the break point Open the debug window to see all the registers. Select Step Into (F11) to run the add line of code Show what values are in the following registers: BL: BH: BX: EBX
Answer #1

Question 1:Assume that you are given values in eax, ebx, ecx. Write an assembly code that does the following: eax = (ecx + edx ) - (eax + ebx)

Solution: This assembly code is an iterpretation of its c counterpart.

************************************************************************************

.LC0:

.string "%d %d %d %d"

.LC1:

.string "%d"

main:

push rbp

mov rbp, rsp

sub rsp, 16

lea rsi, [rbp-16]

lea rcx, [rbp-12]

lea rdx, [rbp-8]

lea rax, [rbp-4]

mov r8, rsi

mov rsi, rax

mov edi, OFFSET FLAT:.LC0

mov eax, 0

call scanf

mov edx, DWORD PTR [rbp-12]

mov eax, DWORD PTR [rbp-16]

lea ecx, [rdx+rax]

mov edx, DWORD PTR [rbp-4]

mov eax, DWORD PTR [rbp-8]

add eax, edx

sub ecx, eax

mov eax, ecx

mov DWORD PTR [rbp-4], eax

mov eax, DWORD PTR [rbp-4]

mov esi, eax

mov edi, OFFSET FLAT:.LC1

mov eax, 0

call printf

mov eax, 0

leave

ret

*******************************************

Question 2:Write a piece of code that copies the number inside al to ch. Example: Assume that Initially eax = 0x15DBCB19. At the end of your code ecx = 0x00001900. Your code must be as efficient as possible

Solution:

************************************

.LC0:

.string "%d"

.LC1:

.string "%d %d"

main:

push rbp

mov rbp, rsp

sub rsp, 16

lea rax, [rbp-4]

mov rsi, rax

mov edi, OFFSET FLAT:.LC0

mov eax, 0

call scanf

mov eax, DWORD PTR [rbp-4]

mov DWORD PTR [rbp-8], eax

lea rdx, [rbp-8]

lea rax, [rbp-4]

mov rsi, rax

mov edi, OFFSET FLAT:.LC1

mov eax, 0

call printf

mov eax, 0

leave

ret

*************************************

Question 4:Implement the following arithmetic expression in assembly language: EBX = var1 - var2 - 9 + var3 // size of variables, 4 bytes each

Solution:

**********************************

main:

push rbp

mov rbp, rsp

mov eax, DWORD PTR [rbp-4]

sub eax, DWORD PTR [rbp-8]

lea edx, [rax-9]

mov eax, DWORD PTR [rbp-12]

add eax, edx

mov DWORD PTR [rbp-16], eax

mov eax, 0

pop rbp

ret

************************************

Question 6:Write a piece of code that implements the following C++ statements d = 'w'; // a ,d,b are of size byte. a = d + 1; b = a + 2;

Solution:

***************************

main:

push rbp

mov rbp, rsp

mov BYTE PTR [rbp-1], 119

movzx eax, BYTE PTR [rbp-1]

add eax, 1

mov BYTE PTR [rbp-2], al

movzx eax, BYTE PTR [rbp-2]

add eax, 2

mov BYTE PTR [rbp-3], al

mov eax, 0

pop rbp

ret

*************************************